A stone was thrown upward at an angle 60° with the horizontal and a vertical speed of 100m/s. What is the actual speed of the stone 10s later in the direction it was thrown?
1
Expert's answer
2020-10-30T04:03:08-0400


At any given time, the velocity of the body will be defined as the geometric sum of these vectors:

V=Vx+Vy

"Vx=V0\times cos\alpha=100\times cos60=100\times0.5=50"

"Vy=V0\times sin\alpha\alpha-g\times t=100\times sin60-10\times10=100\times\frac{\sqrt{3}}{2}-100=-13"

"Vt=\sqrt{(Vx)^2+(Vy)^2}=\sqrt{(50)^2+(-13)^2}=\sqrt{(Vx)^2+(Vy)^2}=\sqrt{2669}=51.66"